# An almost-linear time decoding algorithm for quantum LDPC codes under circuit-level noise

### v1

@@ Abstract
Fault-tolerant quantum computers must be designed in conjunction with classical co-processors that decode quantum error correction measurement information in real-time. In this work, we introduce the belief propagation plus ordered Tanner forest (BP+OTF) algorithm as an almost-linear time decoder for quantum low-density parity-check codes. The OTF post-processing stage removes qubits from the decoding graph until it has a tree-like structure. Provided that the resultant loop-free OTF graph supports a subset of qubits that can generate the syndrome, BP decoding is then guaranteed to converge. To enhance performance under circuit-level noise, we introduce a technique for sparsifying detector error models. This method uses a transfer matrix to map soft information from the full detector graph to the sparsified graph, preserving critical error propagation information from the syndrome extraction circuit. Our BP+OTF implementation first applies standard BP to the full detector graph, followed by BP+OTF post-processing on the sparsified graph. Numerical simulations show that the BP+OTF decoder achieves logical error suppression within an order of magnitude of state-of-the-art inversion-based decoders while maintaining almost-linear runtime complexity across all stages.
@@ I Introduction
Surface code QEC architectures are conceptually simple and can be implemented using local interactions between qubits arranged in a grid. However, a disadvantage of surface codes is that they have poor encoding rate and will require large qubit counts to achieve logical error rates in the teraquop regime (logical failure rate $p_{F}<10^{-12}$) that is thought to be necessary to run algorithms that offer quantum advantage. Quantum low-density parity-check (QLDPC) codes are an alternative QEC protocol to the surface code and are suitable for quantum computing technologies that have high qubit connectivity. Numerical simulations of QLDPC code architectures suggest that they will require approximately 10×10$\times$10 × fewer qubits compared to surface code architectures [5, 6, 7]. However, the design of efficient decoding algorithms for QLDPC codes remains an open problem and a bottleneck to their implementation in experiment.
The belief propagation (BP) decoder is ubiquitous in classical communications, enabling real-time decoding for LDPC protocols underpinning technologies such as 5G, WiFi, and Ethernet  [8]. The primary advantage of BP decoders lies in their speed: they can be executed on distributed hardware and have a linear worst-case runtime complexity of $\mathcal{O}(n)$ with respect to the code block length $n$ [9]. However, when applied to QLDPC codes, BP decoders face significant challenges. These challenges stem from the presence of degenerate quantum errors and the emergence of trapping sets, which can prevent the BP decoder from converging [10, 11]. To address these issues, decoders for QLDPC codes typically augment BP with post-processing routines that are invoked when the main algorithm fails [12, 13, 14, 15, 16, 17, 18]. The most accurate of these two-stage algorithms — ordered statistics decoding (BP+OSD) [13, 14], ambiguity clustering (BP+AC) [16], and localized statistics decoding (BP+LSD) [18] — employ matrix inversion techniques to solve the decoding problem during post-processing. While these inversion-based decoders achieve high error thresholds, they also increase the worst-case runtime complexity to $\mathcal{O}(n^{3})$. This is significantly slower than the $\mathcal{O}(n)$ runtime of standard BP. On the other hand, the closed-branch decoder (BP+CB) significantly reduces the worst-case runtime complexity, but is not as accurate as inversion-based decoders [17].
@@
To benchmark the BP+BP+OTF decoder, we run Monte Carlo simulations under depolarizing circuit-level noise for bivariate bicycle codes, specifically the [[72,12,6]], [[108,8,9]], and [[144,12,12]] instances first investigated in [5]. Our results demonstrate that the BP+BP+OTF decoder achieves error suppression performance that is within an order-of-magnitude of state-of-the art BP+OSD decoders.
Our sparsfication routine for detector error models is general, and has the potential to be useful for other families of decoders beyond BP+BP+OTF. For example, we observe that BP+BP+OSD decoders require considerably fewer BP iterations than is necessary for standalone BP+OSD applied to the full detector graph.
One of the key advantages of BP+BP+OTF is its simplicity: the decoder requires just three applications of a standard BP decoder and a single application of a modification of the well-known Kruskal algorithm for generating the ordered spanning forest. This straightforward design allows for the construction of highly efficient hardware implementations using off-the-shelf application-specific integrated circuits (ASICs). As a result, the BP+OTF decoder is an appealing option for decoding in near-term experiments.
@@ III The Ordered Tanner Forest decoder, an almost-linear time post-processor
In this section, we introduce the ordered Tanner forest (OTF) decoder as an almost-linear post-processor for general QLDPC codes. The OTF decoder is a post-processing algorithm that is invoked if the original round of BP decoding fails, i.e., when the estimated BP error $\mathbf{\hat{e}}_{BP}$ does not satisfy the syndrome equation, $H\cdot\mathbf{\hat{e}}_{BP}\neq\textbf{s}$. Being a post-processor, OTF requires the posterior probabilities coming from a BP decoding stage (or some process able to provide soft reliability information) and is only invoked when BP does not converge. Based on the received soft information, we sort the columns of the parity check matrix from the least reliable to the most reliable [4]. The OTF post-processor then works by deleting columns in the ordered parity check matrix to obtain a reduced parity check matrix, $H_{otf}$, that maps to a Tanner graph that is completely loop free. Note that the resultant spanning tree may contain multiple disconnected trees, so we refer to the returned structure as an ordered Tanner forest (OTF). A product-sum BP decoder is then applied to solve the reduced decoding problem $H_{otf}\cdot\textbf{e}_{otf}=\textbf{s}$. As the $H_{otf}$ matrix has a tree-like structure, the second round of BP is guaranteed to converge if a solution exists, i.e., if $\mathbf{s}$ is linearly dependent on the columns of $H_{otf}$ such that $\textbf{s}\in{\rm\textsc{image}}(H_{otf})$.
@@ IV Sparsifying detector error models
As described in Section II.3, QEC codes are decoded by means of a detector graph that relates circuit fault locations to linear combinations of syndrome measurements. The detector error model enables the propagation of errors through the syndrome extraction circuit to be accounted for during decoding. Due to the richer set of error mechanisms it accounts for, it is typically the case that a detector error model graph will be less sparse than the corresponding CSS parity check matrices $H_{X/Z}$ of the code. This lack of sparsity is detrimental to the success of OTF post-processing, as it increases the probability that the generated OTF matrix will not satisfy the validity condition $\textbf{s}\in{\rm\textsc{image}}(H_{otf})$. See Section III.2 for more details.
@@
Our mapping includes a transfer matrix that allows the error channel associated with each error mechanism in the detector graph to be mapped to an equivalent error channel for the sparsified detector graph. This enables the soft-information output of an initial run of BP applied to the detector model to re-purposed in the sparsified detector graph. The benefit of this is two-fold: first, the sparsified detector graph decoding will account for aspects of error propagation through the circuit, and second, the BP decoding over the OTF graph will be over a non-uniform noise channel, improving its rate of convergence.
@@
We now formally define the sparsified detector error model and its corresponding transfer matrix.
Definition 1 (The sparsified detector matrix)
The detector error model is a $d\times n$ binary matrix $H_{dem}=(\mathbf{h}_{dem}^{1},\mathbf{h}_{dem}^{2},\cdots,\mathbf{h}_{dem}^{n})$, where $\mathbf{h}_{dem}^{i}\in\mathbb{F}_{2}^{d}$ is the binary vector representing the $i^{\rm th}$ column. Let $\Lambda$ be a list of $n_{s}$ indices denoting all columns in $H_{dem}$ that have Hamming weight $|\textbf{h}_{dem}|\leq\gamma$, where $\gamma$ is the maximum Hamming weight of a column in the code’s CSS matrices $H_{X/Z}$. The sparsified detector model, $H_{sdem}$, is a $d\times n_{s}$ sub-matrix of the detector model $H_{sdem}\subseteq H_{dem}$, formed by selecting the low-weight columns of $H_{dem}$ indexed by $\Lambda$. Similarly, the sparsified logical observable matrix $O_{sdem}$ is composed of the subset of columns of $O_{dem}$ indexed by $\Lambda$.
@@
where $\mathbf{e}_{dem}^{i}$ and $\mathbf{e}_{sdem}^{j}$ refer to the circuit faults associated to columns $i$ and $j$ of the detector and spasified models accordingly.
@@
The question as how to best optimise the transfer matrix is an interesting question for future work. For the QEC codes simulated in this work, we found that an exhaustive search method was sufficient. Details of our exhaustive strategy are outlined in Appendix B.
@@ IV.1 Belief propagation decoding using the sparsified detector graph
5:$p(\mathbf{e}_{sdem})\leftarrow$ Map($p(\mathbf{e}_{dem})$$A_{tr}$)
6:$\mathbf{\hat{e}}_{sdem}\leftarrow$ BP_decode($\mathbf{s}$$p(\mathbf{e}_{sdem})$$H_{sdem}$)
@@ V An almost-linear time decoder for QLDPC codes under circuit-level noise
5:$p(\mathbf{e}_{sdem})\leftarrow$ Map($p(\mathbf{e}_{dem})$$A_{tr}$)
6:$(\mathbf{\hat{e}}_{sdem},\ p(\mathbf{e}_{sdem}))\leftarrow$ BP_decode($\mathbf{s}$$p(\mathbf{e}_{sdem})$$H_{sdem}$)
@@
10:$\mathbf{\hat{e}}_{otf}\leftarrow\text{OTF}(H_{sdem},\ p(\mathbf{e}_{sdem}))$
@@
Figure 2: Logical error rate per syndrome extraction round with dependence on the physical error rate for three bivariate bicycle codes under different circuit-level decoding strategies. The left, middle and right plots correspond to bivariate bicycle codes with $l=6,9$ and $12$ respectively, $m=6$, $A=x^{3}+y+y^{2}$ and $B=y^{3}+x+x^{2}$. Each code is simulated over a number of syndrome rounds equal to its distance $\delta$.
@@ VI Results
We numerically benchmark the performance of the BP+BP+OTF decoder via Monte Carlo simulations of the bivariate bicycle codes under a depolarising circuit-level noise model [5]. Precise details of the numerical methods used for this section can be found in Appendix C, whilst a detailed specification of the circuit-level noise model is outlined in Appendix D.
To demonstrate the versatility of the sparsified detector model mapping, we conduct tests with various BP-based decoders. One key advantage of the detector error model sparsification technique is its ability to significantly reduce the total number of iterations required for BP to converge. Below, we outline the decoders we simulate, along with the specific BP iteration counts used for each:
  • Belief propagation (BP): $1000$ iterations over the detector error model graph.
  • Two-stage belief propagation (BP+BP): $30$ iterations over the detector error model and $100$ iterations over the sparsified detector graph.
  • Two-stage BP + ordered Tanner forest (BP+BP+OTF): $30$ iterations over the detector error model, $100$ iterations over sparsified detector graph and $100$ for solving the Tanner forest.
  • Belief propagation + ordered statistics decoding (BP+OSD): $1000$ iterations over the detector error model Tanner graph.
  • Two-stage BP + OSD (BP+BP+OSD): $30$ iterations over the detector error model and $100$ over the sparsified detector error model.
Note that the iteration numbers used for the different decoding algorithms vary significantly. The primary motivation for this variation is to demonstrate that, by employing the proposed detector error model sparsification technique, far fewer iterations of BP are necessary. This reduction in iterations leads to a substantial decrease in runtime over a straight-forward application of BP for $N_{BP}$ iterations.
@@ VI.1 Bivariate bicycle codes
Figure 2 shows the results of Monte Carlo decoding simulations for families of bivariate bicycle codes. Decoding simulations were run for the $[[72,12,6]]$, $[[108,8,10]]$, and $[[144,12,12]]$ bivariate bicycle codes first introduced in [5]. Notably, the proposed BP+BP protocol significantly enhances performance compared to running BP directly on the detector error model Tanner graph across all the codes. Specifically, we observe nearly an order-of-magnitude reduction in the logical error rate. Importantly, this improvement is achieved with approximately 90% fewer overall BP iterations (1000 vs. 130), resulting in a significantly reduced runtime. However, the results also indicate that simply increasing the code size does not yield an improvement in the logical error rate when using BP alone. Therefore, post-processing techniques are still necessary to achieve the desired threshold-like behavior for the bivariate bicycle codes, as anticipated.
Figure 2 illustrates that the performance of BP+OSD is effectively matched by the BP+BP+OSD approach evaluated here. The key point is that this performance parity is achieved with approximately 90% fewer BP iterations overall. Similarly, the proposed BP+BP+OTF achieves a logical error probability within an order of magnitude of BP+OSD, again using around 90% fewer BP iterations. Furthermore, the OTF post-processor has a proven almost-linear runtime complexity, ensuring that the entire decoding process is extremely fast.
It is also noteworthy that the error curve for BP+BP+OTF steadily diverges from those of BP+OSD and BP+BP+OSD as the code distance increases. This divergence may be due to the larger size of the OTF, which might require additional BP iterations to fully converge. In future work, we plan to investigate the number of BP rounds necessary for BP+BP+OTF to match the performance of running BP+OSD on the detector error model.

### v2

@@ Abstract
Fault-tolerant quantum computers must be designed in conjunction with classical co-processors that decode quantum error correction measurement information in real-time. In this work, we introduce the belief propagation plus ordered Tanner forest (BP+OTF) algorithm as an almost-linear time decoder for quantum low-density parity-check codes. The OTF post-processing stage removes qubits from the decoding graph until it has a tree-like structure. Provided that the resultant loop-free OTF graph supports a subset of qubits that can generate the syndrome, BP decoding is then guaranteed to converge. To enhance performance under circuit-level noise, we introduce a technique for sparsifying detector error models. This method uses a transfer matrix to map soft information from the full detector graph to the sparsified graph, preserving critical error propagation information from the syndrome extraction circuit. Our BP+OTF implementation first applies standard BP to the full detector graph, followed by BP+OTF post-processing on the sparsified graph. Numerical simulations show that the BP+OTF decoder achieves similar logical error suppression compared to state-of-the-art inversion-based and matching decoders for bivariate bicycle and surface codes, respectively, while maintaining almost-linear runtime complexity across all stages.
@@ I Introduction
Surface code QEC architectures are conceptually simple and can be implemented using local interactions between qubits arranged in a grid. However, a disadvantage of surface codes is that they have poor encoding rate and will require large qubit counts to achieve logical error rates in the teraquop regime (logical failure rate $p_{L}<10^{-12}$) that is thought to be necessary to run algorithms that offer quantum advantage. Quantum low-density parity-check (QLDPC) codes are an alternative QEC protocol to the surface code and are suitable for quantum computing technologies that have high qubit connectivity. Numerical simulations of QLDPC code architectures suggest that they will require approximately $10\times$ fewer qubits compared to surface code architectures [5, 6, 7]. However, the design of efficient decoding algorithms for QLDPC codes remains an open problem and a bottleneck to their implementation in experiment.
The belief propagation (BP) decoder is ubiquitous in classical communications, enabling real-time decoding for LDPC protocols underpinning technologies such as 5G, WiFi, and Ethernet  [8]. The primary advantage of BP decoders lies in their speed: they can be executed on distributed hardware and have a linear worst-case runtime complexity of $\mathcal{O}(n)$ with respect to the code block length $n$ [9]. However, when applied to QLDPC codes, BP decoders face significant challenges. These challenges stem from the presence of degenerate quantum errors and the emergence of trapping sets, which can prevent the BP decoder from converging [10, 11]. To address these issues, decoders for QLDPC codes typically augment BP with post-processing routines that are invoked when the main algorithm fails [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The most accurate of these two-stage algorithms — ordered statistics decoding (BP+OSD) [13, 14], ambiguity clustering (BP+AC) [16], and localized statistics decoding (BP+LSD) [18] — employ matrix inversion techniques to solve the decoding problem during post-processing. While these inversion-based decoders achieve high error thresholds, they also increase the worst-case runtime complexity to $\mathcal{O}(n^{3})$. This is significantly slower than the $\mathcal{O}(n)$ runtime of standard BP.
@@
To benchmark the BP+BP+OTF decoder, we run Monte Carlo simulations under depolarising circuit-level noise for bivariate bicycle codes [5] as well as for surface codes. Our results demonstrate that the BP+BP+OTF decoder achieves comparable error suppression to state-of-the-art decoders such as BP+OSD and minimum-weight perfect-matching. Notably, our implementation runs more than $10\times$ faster compared to the BP+OSD implementation from [26] when decoding bivariate bicycle codes under certain noise regimes. Furthermore, our results show that BP+BP+OTF can reliably decode surface codes using a very small number of iterations of the minimum-sum algorithm.
One of the key advantages of BP+BP+OTF is its simplicity: the decoder requires just three applications of a standard BP decoder and a single application of a modification of the well-known Kruskal algorithm for generating the ordered spanning forest. This straightforward design allows for the construction of highly efficient hardware implementations using off-the-shelf application-specific integrated circuits (ASICs). As a result, the BP+OTF decoder is an appealing option for real-time decoding.
@@ III The Ordered Tanner Forest decoder, an almost-linear time post-processor
In this section, we introduce the ordered Tanner forest (OTF) decoder as an almost-linear post-processor for general QLDPC codes. The OTF decoder is a post-processing algorithm that is called if the original round of BP decoding fails, i.e., when the estimated BP error $\mathbf{\hat{e}}_{BP}$ does not satisfy the syndrome equation, $H\cdot\mathbf{\hat{e}}_{BP}\neq\textbf{s}$. Being a post-processor, OTF requires the posterior probabilities coming from a BP decoding stage (or some process able to provide soft reliability information) and is only invoked when BP does not converge. Based on the received soft information, we sort the columns of the parity check matrix from the least reliable to the most reliable [4]. The OTF post-processor then works by deleting columns in the ordered parity check matrix to obtain a reduced parity check matrix, $H_{otf}$, that maps to a Tanner graph that is completely loop free. Note that the resultant spanning tree may contain multiple disconnected trees, so we refer to the returned structure as an ordered Tanner forest (OTF). A product-sum BP decoder is then applied to solve the reduced decoding problem $H_{otf}\cdot\textbf{e}_{otf}=\textbf{s}$. As the $H_{otf}$ matrix has a tree-like structure, the second round of BP is guaranteed to converge if a solution exists, i.e., if $\mathbf{s}$ is linearly dependent on the columns of $H_{otf}$ such that $\textbf{s}\in{\rm\textsc{image}}(H_{otf})$.
@@ IV Sparsifying detector error models
As described in Section II.3, QEC codes are decoded by means of a detector graph that relates circuit fault locations to linear combinations of syndrome measurements. The detector error model enables the propagation of errors through the syndrome extraction circuit to be accounted for during decoding. Due to the richer set of error mechanisms it considers, it is typically the case that a detector error model graph will be less sparse than the corresponding CSS parity check matrices $H_{X/Z}$ of the code. This lack of sparsity is detrimental to the success of OTF post-processing, as it increases the probability that the generated OTF matrix will not satisfy the validity condition $\textbf{s}\in{\rm\textsc{image}}(H_{otf})$. See Section III.2 for more details.
@@
Our mapping includes a transfer matrix that allows the error channel associated with each error mechanism in the detector graph to be mapped to an equivalent error channel for the sparsified detector graph. This enables the soft-information output of an initial run of BP applied to the detector model to be re-purposed in the sparsified detector graph. The benefit of this is two-fold: first, the sparsified detector graph decoding will account for aspects of error propagation through the circuit, and second, the BP decoding over the OTF graph will be over a non-uniform noise channel, improving its rate of convergence.
@@
We now formally define the sparsified detector error model and its associated transfer matrices.
Definition 1 (Sparsified Detector Matrix)
Let the detector error model be represented by a binary matrix $H_{\mathrm{dem}}\in\mathbb{F}_{2}^{d\times n}$, where $H_{\mathrm{dem}}=(\mathbf{h}_{\mathrm{dem}}^{1},\mathbf{h}_{\mathrm{dem}}^{2},\ldots,\mathbf{h}_{\mathrm{dem}}^{n})$ and each column $\mathbf{h}_{\mathrm{dem}}^{i}\in\mathbb{F}_{2}^{d}$. The sparsified detector matrix, denoted $H_{\mathrm{sdem}}\in\mathbb{F}_{2}^{d\times n_{s}}$ with $n_{s}<n$, is defined such that
${\rm\textsc{image}}(H_{\mathrm{sdem}})={\rm\textsc{image}}(H_{\mathrm{dem}}),$
i.e., the columns of $H_{\mathrm{sdem}}$ form a generating set for the column space of $H_{\mathrm{dem}}$. Similarly, we define the sparsified logical observable matrix $O_{\mathrm{sdem}}$, whose columns lie in the image of the original logical observable matrix $O_{\mathrm{dem}}$.
Note that the selection of the sparsified detector matrix is not unique. In this work, we use the following sparsified models:
  • For surface codes, the standard circuit-level matching graph is used as the sparsified detector model. This is obtained by decomposing hyperedges in the full detector graph into matchable edges [33].
  • For bivariate bicycle codes, a quasi-phenomenological model is used as the sparsified detector graph. This is obtained by removing the two-qubit gate errors from the circuit-level noise model.
We observe that the number of columns in the sparsified detector models is typically $\approx 30\%$ of the original DEM. This shows that the models are not only sparser, but also much smaller.
@@
where $\mathbf{e}_{dem}^{i}$ and $\mathbf{e}_{sdem}^{j}$ refer to the circuit faults associated to columns $i$ and $j$ of the detector and sparsified models accordingly.
@@
The question as how to best optimise the sparsified detector model and the transfer matrix is an interesting question for future work. For the QEC codes simulated in this work, we found that an exhaustive search method was sufficient. Details of our exhaustive strategy are outlined in Appendix B.
@@ IV.1 Belief propagation decoding using the sparsified detector graph
[…]
@@ V An almost-linear time decoder for QLDPC codes under circuit-level noise
[…]
@@
[…]
@@
[…]
Figure 2: Logical error rate per syndrome extraction round as a function of the physical error rate for bivariate bicycle codes. Each code is simulated over a number of syndrome rounds equal to its distance. The shading highlights the region of estimated probabilities where the likelihood ratio is within a factor of $1,000$; similar to a confidence interval.
@@ VI Results
We numerically benchmark the performance of the BP+BP+OTF decoder via Monte Carlo simulations of the bivariate bicycle codes and rotated surface codes under a depolarising circuit-level noise model [5]. Precise details of the numerical methods used for this section can be found in Appendix C, whilst a detailed specification of the circuit-level noise model is outlined in Appendix D.
@@ VI.1 Bivariate Bicycle Codes
Figure 2 shows the logical error rate per round for families of bivariate bicycle codes decoded using BP+OSD and BP+BP+OTF. Simulations were performed on the $[[72,12,6]]$, $[[108,8,10]]$, $[[144,12,12]]$, and $[[288,12,18]]$ bivariate bicycle codes first introduced in [5]. The decoder setup is as follows:
  • BP+OSD-0: We set the maximum number of iterations for the minimum-sum belief propagation decoder to 10,00010{,}000, following [5]. As noted in [16], using a large iteration count improves speed, as premature termination means the decoder relies more on OSD-postprocessing. The ‘0’ in OSD-0 indicates that no higher-order post-processing is applied find solutions that improve upon the initial output [14].
  • BP+BP+OTF: Oscillations in the soft output of BP can emerge after a sufficient number of iterations [19], which may result in OTF graphs that fail to span the observed syndrome. To address this, we adopt an ensemble decoding strategy [36, 22], where each member of the ensemble uses a different number of iterations in the first BP stage. This improves robustness without affecting overall runtime, as all decoders run in parallel  [^3] For the $[[72,12,6]]$ code, which exhibits limited degeneracy, we do not use the ensemble approach. Instead, we run a single BP+BP+OTF decoder with $300$ iterations on the full detector graph and $113$ iterations in the second BP stage and on the OTF graph.. The ensemble consists of $22$ BP+BP+OTF decoders, with first-stage BP iteration counts ranging from $17$ to $391$ in steps of $17$. The subsequent BP rounds on both the sparsified and OTF graphs are capped at $113$ iterations. All BP stages use the minimum-sum variant.
[…]
(a) physical error rate @ $p=0.003$.
[…]
(b) physical error rate @ $p=0.004$.
Figure 3: Mean decoding time per round for BP+OSD-0 and BP+BP+OTF when decoding bivariate bicycle codes. For the BP+BP+OTF ensemble we record the time of the first instance that finishes decoding. a) Decoding at physical error rate $p=0.003$. b) Decoding at physical error rate $p=0.004$.
The results demonstrate that the proposed decoder matches the performance of BP+OSD-0 when decoding bivariate bicycle codes. Crucially, this is achieved with an almost-linear runtime, requiring at most $617$ BP iterations across the three stages of the decoder.
To evaluate the speedup provided by our implementation, we also performed timing benchmarks. Full details of these tests are provided in Appendix C. Figure 3 reports the mean decoding time per syndrome extraction round for both decoders at two physical error rates, $p=0.003$ and $p=0.004$ (see Appendix E for detailed statistics). The results clearly show that our decoder consistently outperforms the standard BP+OSD decoder, achieving speedups greater than a factor of $10$ across all scenarios.
This performance gap should be interpreted with care. The BP implementation used from [26] executes serially, even under a parallel schedule. As a result, the actual complexity of our current implementation is quadratic. A fully parallel implementation would yield significantly greater speedups compared to BP+OSD. Nonetheless, these results already highlight that the BP+BP+OTF decoder exhibits much better scaling behaviour, particularly for codes with higher distances and larger detector error models.
Another important aspect of our analysis concerns the amount of decoding information lost when the detector error model (DEM) is sparsified. To investigate this, we conduct a series of numerical experiments using BP+OSD-0 in different configurations, with detailed results presented in Appendix F.
First, we compare the performance of BP+OSD-0 applied directly to the full DEM with the same decoder applied to the sparsified model, both using a maximum of 10,000 BP iterations. We observe a significant degradation in performance, with logical error rates worsening by several orders of magnitude, and no improvement in decoding time. This indicates that completely ignoring the structure of the full DEM is not an effective strategy.
We then evaluate the BP+BP+OSD-0 decoder, where the first BP stage runs on the full DEM for 1,000 iterations and the second stage runs on the sparsified model for 9,000 iterations, followed by OSD-0 on the sparsified model. In this setup, 90 percent of the BP iterations and all post-processing are performed on the sparsified graph. Remarkably, this hybrid approach matches the logical error rate of BP+OSD-0 executed entirely on the full DEM, while achieving a threefold reduction in average decoding time.
These results support the effectiveness of our sparsification technique. Running a limited number of BP iterations on the full DEM is sufficient to extract the essential soft information required for decoding. Since the sparsified model is also smaller, other decoders that perform well under phenomenological noise may benefit from this approach, enabling them to operate effectively under circuit-level noise with reduced dependence on the full DEM.
VI.2 Surface codes
[…]
(a) BP+BP+OTF versus MWPM.
[…]
(b) BP+BP+OTF versus BP+OSD-0.
Figure 4: Logical error rate per syndrome extraction round as a function of the physical error rate for rotated surface codes. Each code is simulated over a number of syndrome rounds equal to its distance using Stim’s surface_code:rotated_memory_z experiment. The shading highlights the region of estimated probabilities where the likelihood ratio is within a factor of $1,000$; similar to a confidence interval.
Figure 4 presents the logical error rates per round for rotated surface codes with distances d=5,9,13,17d={5,9,13,17}, decoded using minimum-weight perfect matching (MWPM), BP+OSD-0, and BP+BP+OTF. For the BP+OSD-0 decoder, we use 70 minimum-sum BP iterations, while the BP+BP+OTF decoder employs 6, 51, and 50 maximum iterations for its three decoding stages, respectively. As shown in the figure, BP+BP+OTF achieves logical error suppression comparable to MWPM (Figure 4a), and significantly outperforms BP+OSD-0 (Figure 4b). This result is particularly notable, as BP+OSD-0 is known to slightly outperform belief-find [18], a decoder with almost-linear runtime complexity. These findings highlight the strong performance of BP+BP+OTF in both accuracy and efficiency.
A distinctive feature of the BP+BP+OTF decoder in the context of surface codes is its guaranteed convergence. This property arises from the fact that the OTF post-processor operates on a matrix with a matching structure. In the case of surface codes without periodic boundaries, an additional virtual node must be included in the matrix to enable OTF decoding (see Appendix G). This results in an augmented matrix with $d+1$ rows, and each column has weight 2. Since a tree-like graph must satisfy |V||E|=1|V|-|E|=1, where $|V|$ and $|E|$ are the numbers of vertices and edges, respectively, we can deduce the number of columns in the OTF matrix. Letting $n_{\text{otf}}$ denote the number of such columns, the relationship becomes $(d+1+n_{\text{otf}})-2n_{\text{otf}}=1$, which implies $n_{\text{otf}}=d$. The final BP step is then performed on a full-rank $d\times d$ matrix, excluding the virtual node, ensuring that the syndrome is always in the image of $H_{\text{otf}}$. This guarantees convergence and makes BP+BP+OTF particularly attractive for codes with matching structure [4].
We do not explicitly benchmark the runtime of our BP+BP+OTF implementation against pymatching [38], since pymatching is highly optimised for CPU execution, while our BP implementation is serial and not performance-tuned. Nonetheless, it is noteworthy that BP+BP+OTF is capable of decoding a $d=17$ surface code using a total of only 107 BP iterations. This suggests that a hardware-accelerated implementation, for example on an FPGA, could achieve decoding speeds and meet the requirements for real-time operation [23].